### v1

# Reinforcement Learning vs. Distillation: Understanding Accuracy and Capability in LLM Reasoning

Abstract
Recent studies have shown that reinforcement learning with verifiable rewards (RLVR) enhances overall accuracy but fails to improve capability, while distillation can improve both. In this paper, we investigate the mechanisms behind these phenomena. First, we demonstrate that RLVR does not improve capability because it focuses on improving the accuracy of the less-difficult questions to the detriment of the accuracy of the most difficult questions, thereby leading to no improvement in capability. Second, we find that RLVR does not merely increase the success probability for the less difficult questions, but in our small model settings produces quality responses that were absent in its output distribution before training. In addition, we show these responses are neither noticeably longer nor feature more reflection-related keywords, underscoring the need for more reliable indicators of response quality. Third, we show that while distillation reliably improves accuracy by learning strong reasoning patterns, it only improves capability when new knowledge is introduced. Moreover, when distilling only with reasoning patterns and no new knowledge, the accuracy of the less-difficult questions improves to the detriment of the most difficult questions, similar to RLVR. Together, these findings offer a clearer understanding of how RLVR and distillation shape reasoning behavior in language models. [^1] Code available at https://github.com/minwukim/RLvsDistillation
Reinforcement Learning vs. Distillation:
Understanding Accuracy and Capability in LLM Reasoning
Minwu Kim*  Anubhav Shrestha*  Safal Shrestha  Aadim Nepal  Keith Ross New York University Abu Dhabi
11footnotetext: Equal contribution.footnotetext: Correspondence: keithwross@nyu.edu
1 Introduction
Large language models (LLMs) have made rapid progress in complex domains such as mathematics and programming. A key development is the emergence of reasoning models OpenAI (2024); DeepSeek-AI (2025); MoonshotAI (2025), which outperform conventional LLMs by employing more advanced reasoning strategies. Instead of relying on linear chains of thought (CoTs) Wei et al. (2022), these models exhibit non-linear behaviors such as subgoal formulation, verification, backtracking, and backward chaining Gandhi et al. (2025); Xiang et al. (2025).
A central technique behind recent reasoning models is reinforcement learning with verifiable rewards (RLVR) Lambert et al. (2024); DeepSeek-AI (2025). RLVR fine-tunes pre-trained LLMs using rewards based on whether the model’s output matches a ground-truth solution. Without explicit supervision, the model learns complex reasoning behaviors during training, making RLVR an effective fine-tuning method for reasoning tasks. In addition, distillation with responses from stronger reasoning models also provide strong performance, demonstrating that reasoning ability can be effectively transferred through supervised learning Min et al. (2024); DeepSeek-AI (2025); Muennighoff et al. (2025); Ye et al. (2025).
It has been recently observed that RLVR improves accuracy—the probability that a correct answer is generated, but does not improve capability-whether the correct answers exists within the model’s output distribution Yue et al. (2025); Shao et al. (2024). It has been further observed that distillation improves both accuracy and capability Yue et al. (2025). This prior research points out important “whats” for RLVR and teacher distillation, and provides an interesting contrast among the two techniques. In this paper, we take a deeper dive into these issues by examining not only the “whats” but also “why” these phenomena occur.
Carrying out experiments with two models, Qwen2.5-1.5B-Math Yang et al. (2024) and Qwen2.5-3B Hui et al. (2024), we first demonstrate that RLVR does not improve capability because RLVR focuses on improving the accuracy of the less-difficult questions to the detriment of the accuracy of the most difficult questions, explaining why capability does not improve and can even decrease. We argue that this "sacrificing-difficult-problems" phenomenon is a direct consequence of the underlying policy-gradient algorithm in GRPO Shao et al. (2024). We further find that RLVR does not merely increase the success probability for the less difficult questions, but in our small model settings produces responses that are more direct with fewer keywords. In addition, we find these responses are neither noticeably longer nor richer in reflection-related keywords (e.g., “wait”, “alternatively”), underscoring the need for more reliable indicators of reasoning quality.
Finally we turn our attention to teacher distillation. A teacher model’s responses carry two main types of information: (1) its reasoning patterns and (2) its domain knowledge. To disentangle their effects, we compare three models: the base model; the publicly released DeepSeek reasoning model, which was reportedly distilled 800k responses from DeepSeek-R1 and likely incorporates substantial new knowledge; and our own distilled model, trained only on teacher responses for questions whose correct answers already existed in the base model’s output distribution. We find that although both distilled models significantly improve accuracy, only the DeepSeek model shows a meaningful increase in capability. Thus the reason why teacher distillation improves both accuracy and capability is not because it brings new reasoning patterns to the distilled model but instead brings new knowledge, knowledge which was not present in the original model. Comparing RLVR and teacher distillation is thus an apples-to-oranges comparison, in that RLVR only uses the resources it has on hand, whereas teacher distillation augments the model with new resources. We also make one other interesting observation: when distilling only with reasoning patterns and no new knowledge, the accuracy of the less-difficult questions improves to the detriment of the most difficult questions, similar to RLVR.
Taken together, our findings provide a clearer picture of behind the "whys" of RLVR-trained and teacher-distilled reasoning models and help to contrast the two approaches.
2 Related Work
2.1 Reinforcement Learning with Verifiable Rewards (RLVR)
RLVR has emerged as an effective method for training LLMs to solve complex reasoning tasks Lambert et al. (2024). Notable demonstrations include OpenAI’s o1 OpenAI (2024), DeepSeek-R1 DeepSeek-AI (2025), and Kimi k1.5 MoonshotAI (2025), which use RLVR to produce long CoTs and achieve substantial gains across various benchmarks. Follow-up studies have replicated these successes across different model sizes Gandhi et al. (2025); Hu et al. (2025); Liu et al. (2025); Xu et al. (2025); Yeo et al. (2025); Zeng et al. (2025), as well as in diverse domains such as medical and synthetic reasoning tasks Pan et al. (2025); Xie et al. (2025); Zhang et al. (2025). In parallel, various algorithmic variants of RLVR have been proposed to improve model performance, data efficiency, and computational cost Shao et al. (2024); Fatemi et al. (2025); Liu et al. (2025); Wang et al. (2025); Yu et al. (2025); Zuo et al. (2025). However, most prior work has focused on improving the effectiveness of RLVR itself. In contrast, our work aims to deepen the understanding of how RLVR changes model behavior. By analyzing its effects through the dual perspectives of accuracy and capability, we clarify the mechanisms by which RLVR shapes reasoning in LLMs.
2.2 Distillation for Reasoning in LLMs
Distilling high-quality chain-of-thought (CoT) data has also proven to be an effective approach for improving the reasoning ability of LLMs. Various efforts have focused on collecting such data. In some cases, high-quality reasoning traces are generated by prompting large language models Zelikman et al. (2022); Yu et al. (2024); in others, human annotators construct answers that reflect complex, non-linear reasoning patterns Ye et al. (2025); Qin et al. (2024); Xiang et al. (2025). More recently, a popular strategy involves distilling responses from RLVR-trained reasoning models into student models, often resulting in significant performance improvements Huang et al. (2024); Min et al. (2024); Muennighoff et al. (2025). In parallel, several studies have explored which qualitative properties of CoT demonstrations—such as problem difficulty or structural format—most strongly influence the effectiveness of distillation Li et al. (2025); Sun et al. (2025). What sets our work apart is a fine-grained analysis of the distillation data itself. We disentangle the effects of reasoning style and domain knowledge, examining how each influences both model accuracy and capability. This perspective provides deeper insight into how reasoning-focused distillation shapes model behavior.
3 Accuracy and Capability
3.1 Formal Definitions
We evaluate models along two dimensions: accuracy and capability. Informally, accuracy measures how likely a model is to generate a correct answer in a single attempt, while capability measures whether a correct answer exists within the distribution of responses the model can generate.
Formally, we define accuracy and capability with respect to given model $M$ and evaluation dataset $\mathcal{D}=\{1,\dots,N\}$ of $N$ questions. Let $p_{i}^{M}$ denote the probability that model $M$ successfully solves question $i$ in a single attempt. Note that this can be obtained by sampling model $M$ for $n$ times on question $i$, computing the fraction of correct responses, and taking the limit as $n\to\infty$. In theory, an LLM using softmax sampling assigns non-zero probability to all valid outputs, so any answer could eventually be produced. To make capability practically meaningful, we consider a question $i$ to be in-distribution for model $M$ if $p_{i}^{M}>\epsilon$, where $\epsilon$ is a small threshold (typically $10^{-2}$ to $10^{-3}$).
To evaluate performance under multiple attempts, let $p_{i,k}^{M}$ denote the probability that model $M$ solves question $i$ at least once across $k$ independent attempts. This probability satisfies
$p_{i,k}^{M}=1-(1-p_{i}^{M})^{k}.$
With these definitions in place, we define the model’s accuracy as the average success rate over the entire dataset:
$\mathrm{Acc}(M)=\frac{1}{N}\sum_{i\in\mathcal{D}}p_{i}^{M}.$
We define the model’s pass@k capability as the average success probability over $\mathcal{D}$ given $k$ passes per question:
$\mathrm{Cap}_{k}(M)=\frac{1}{N}\sum_{i\in\mathcal{D}}p_{i,k}^{M}=\frac{1}{N}\sum_{i\in\mathcal{D}}\left(1-(1-p_{i}^{M})^{k}\right)$
It is important to note that if model $M^{\prime}$ has higher accuracy than model $M$ ($p_{i}^{M^{\prime}}>p_{i}^{M}$) for a specific question, then it will also have higher pass@k capability for that question. However, this relationship does not always hold taking into account the entire dataset. In fact, as shown in Appendix A.1, it is possible for $\mathrm{Acc}(M^{\prime})>\mathrm{Acc}(M)$ while $\mathrm{Cap}_{k}(M^{\prime})<\mathrm{Cap}_{k}(M)$.
3.2 Estimating Accuracy and Capability
In practice, it is infeasible to compute the exact accuracy and capability of a model, as this would require a prohibitively large number of samples per question. Instead, we estimate these quantities empirically using a finite number of samples $k$. Let $X_{i,k}$ be the number of correct responses among $k$ samples for question $i$.
We estimate accuracy as:
$\mathrm{Acc}(M)\approx\frac{1}{N}\sum_{i\in\mathcal{D}}\frac{X_{i,k}}{k}$
We estimate pass@k capability as:
$\mathrm{Cap}_{k}(M)\approx\frac{1}{N}\sum_{i\in\mathcal{D}}1(X_{i,k}>0)$
These estimators are unbiased. Throughout this work, we report results using these estimators, typically with $k=256$. We also consider a question $i$ to be out-of-distribution if $X_{i,256}=0$, that is, none of the 256 responses to question $i$ are correct. Under this definition, we can say with 95% confidence that $p_{i}^{M}<1-(0.05)^{1/256}\approx 0.012$, that is, question $i$ is truly out-of-distribution under the threshold $\epsilon=0.012$.
4 Why Doesn’t RLVR Improve Capability?
In prior work, it was shown empirically from various models that RLVR can provide significant improvements in accuracy but in doing so it may not improve pass@$k$ capability when $k$ is chosen to be sufficiently large Yue et al. (2025). In this section, we build on that observation and examine the phenomenon in greater depth. Our analysis also lays the foundation for the experiments that follow. Specifically, we aim to answer the following question: Why does RLVR improve accuracy while leaving capability unchanged or even degraded?
4.1 Capability Analysis
We first performed pass@$k$ experiments similar to those by Yue et al., confirming that RLVR improves accuracy but not capability (Fig. 7). Specifically, we evaluated two base models—Qwen2.5-1.5B-Math Yang et al. (2024) and Qwen2.5-3B Hui et al. (2024)—along with their corresponding RLVR-trained versions. The full training details and pass@$k$ results are provided in Appendix A.9 and A.2, respectively. All models were trained on the MATH training set, which contains 7,500 questions, and evaluated on the MATH 500 test set with 500 problems. These same datasets are also used for all subsequent experiments in this section.
We report here results on the 1.5B model evaluated on the test set due to space constraints. However, the same pattern holds consistently across both the train and test sets, and across both model sizes (1.5B and 3B). Full results are provided in Appendix A.3. For clarity, we refer to the original 1.5B model as the base model and the RLVR-trained version as the RL model.
4.2 A Deeper-Dive: Analysis Based on Question Difficulty
Figure 1: Change in success rates after RLVR training on MATH 500 test set, grouped by the base model’s per-question success rate measured over 256 responses. The bar height represents the absolute difference (%) between the RL and base models within each bin.
To better understand the accuracy-capability dynamics of RLVR, we perform a fine-grained analysis based on question-level success rates. Specifically, we perform the following experiment: using the 256 responses generated by the base model for each question in the training and test sets, we compute a per-question success rate. We then group the questions into bins based on these success rates: [0], [1–4], [5–16], [17–64], [65–128], and [129-256]. For each bin, we identify the corresponding questions and retrieve the RL model’s responses to those same questions. Finally, we compute the average success rate for both the base and RL models within each bin, and plot the difference between them. The results are shown in Figure 1.
Figure 2: Success rate transition matrix showing redistribution of questions from base model to RL model on MATH 500 test set.
We observe two consistent patterns. (1) For questions where the base model already has a moderately high success rate, the RL model shows substantial improvement. For example, in the training set, the [65–128] bin shows a 36.6% gain in average success rate relative to 256 responses, and the [17–64] bin shows a 24.7% gain. (2) In contrast, for questions where the base model has zero or near-zero success rate, the improvement is negligible. For instance, the [0] bin in the training set improves by only 0.5% points, and the [1–4] bin by just 0.6% points. These patterns are consistent across both the train and test sets and across both model sizes (1.5B and 3B), as shown in Appendix A.3.
Furthermore, to better understand the pattern observed in success rate improvements, we visualize how individual questions move across success rate bins before and after RLVR training. Figure 2 presents this transition in test set as a matrix. Here, each row corresponds to a success-rate bin based on the base model’s performance, and each column corresponds to the same bin based on the RL model’s performance. Each cell shows the percentage (and count) of questions that started in a specific base model bin and end up in a particular RL model bin after training.
Notably, we observe two clear trends. (1) Questions that were already in high-success bins tend to stay in the same bin or move to even higher bins after RLVR training. For instance, in the [65–128] bin, 15.4% remain in place and striking 83.3% move to higher [129-256] bin—while only 1.3% (1 question out of 78) drop to lower ones. Similar upward shifts are seen in the [17–64] bin. (2) In contrast, questions that started in low-success bins, especially those with near-zero success rates, are more likely to stay in the same bin or shift further downward. For example, in the [1-4] bin, a substantial 61.1% remain in the same bin while 16.7% drop down to the [0] bin, and a similar pattern appears in the [5–16] range where 67.4% of the questions stay in the same bin or regress. This shows a clearer picture of RLVR failing to help with previously unsolved questions and can even increase their number. Many questions with a small chance of being answered correctly end up never being answered correctly at all after training. Again, these trends are consistent across both the train and test sets and across both 1.5B and 3B models, as discussed in Appendix A.3.
To understand this behavior, we need to examine the internal dynamics of RLVR training. Take GRPO algorithm as an example. At each iteration, the model generates multiple responses (e.g., 8) per question in a batch. Each response is evaluated for correctness, and parameter updates are applied accordingly. If all 8 responses for a given question are incorrect, that question has no influence on the update. In contrast, for questions with a mix of correct and incorrect responses, the model parameters are nudged to increase the likelihood of generating correct answers and decrease incorrect ones. Now consider an early training step where the batch includes one difficult question (e.g., in bin [1–4]) and several easier ones (e.g., in bin [65–128]). With high probability, the model will generate only incorrect responses for the difficult question, while producing at least some correct answers for the easier ones. As a result, the parameter update will be guided entirely by the easier questions only. This can lead to the model assigning even lower probability to the correct answers for the difficult question, especially if model capacity is limited. This selective reinforcement continues throughout training, explaining why questions that initially had a small chance of being answered correctly may become even less likely to be solved after training. To confirm this selective reinforcement effect, we also analyze the entropy of the model’s output distributions—measured over 256 sampled responses per question—across different difficulty levels. We observe that entropy consistently decreases after RLVR training. Full results are provided in Appendix A.4.
To summarize, these results suggest to the following insight: RLVR improves accuracy but not capability because RLVR focuses on improving the accuracy of the less-difficult questions to the detriment of the accuracy of the most difficult questions.
5 Comparison of Model Responses Before & After RLVR
RLVR causes the model to increase the probability of generating correct answers it could already generate, but does not enable it to solve previously unsolvable questions. This naturally leads us to the following question: Can we replicate RLVR’s effect by directly nudging the model toward its own correct responses—that is, through self-distillation? And if not, why?
5.1 Self-Distillation Experiments
Setup Train Acc (%) Test Acc (%)
1) Base 64.0 62.6
2) Base $\leftarrow$ Base 74.7 (+10.7) 63.4 (+0.8)
3) RL 80.9 74.8
4) RL    $\leftarrow$ RL 84.4 (+3.5) 74.4 (-0.4)
5) Base $\leftarrow$ RL 80.5 (+16.5) 74.2 (+11.6)
Table 1: Self-distillation results for the Qwen2.5-1.5B-Math base model and its RLVR-trained variant. The notation $A$leftarrow $B$ indicates the student model $A$ is trained on responses from the teacher model $B$. Values in parentheses show gains over the student model with distillation.
To explore this question, we perform self-distillation with rejection sampling, following approaches similar to STaR Zelikman et al. (2022) and ReST Gulcehre et al. (2023)—that is, we conduct SFT on the model’s own correct responses. Recall that, we collected 256 responses per training question from the base model in Section 4. From these, we select up to 8 correct responses for each question (using all available correct ones if fewer than 8 exist) and conduct SFT using this filtered dataset. Additionally, to test whether self-distillation can also further improve the RLVR-trained model, we apply the same self-distillation process to the RL model. We conduct this experiment for both Qwen2.5-1.5B-Math and Qwen2.5-3B, but present only the 1.5B results here due to space constraints. The 3B results show similar trends and are provided in Appendix A.5.
We find that self-distillation fails to replicate the effect of RLVR. As shown in Table 1 (lines 1-2), distilling the base model using its own correct responses yields a train accuracy of 74.7%, a 10.7-point increase. However, test accuracy rises only modestly to 63.4%, a 0.8-point gain over the base model’s 62.6% and significantly below the RL model’s 74.8%. Similarly, from lines 3-4 we observe self-distillation of the RL model leads to no gain on the test set (74.8% to 74.4%), despite notable rises in training accuracy (80.9% to 84.4%).
These results suggest that, unlike RLVR, self-distillation tends to overfit to the training set and fails to promote more generalizable reasoning behavior.
We take a step further and conduct another self-distillation experiment. Prior work has shown that distilling quality responses from a teacher model can effectively improve the performance of student model  Huang et al. (2024); Min et al. (2024); Muennighoff et al. (2025). In other words, responses that lead to large accuracy gains through distillation can be seen as having high quality. Based on this idea, we perform an additional experiment: distilling the RL model’s responses into the base model.
Interestingly, the base model shows a significant performance gain that is not limited to the training set. As shown in Table 1 (line 5), its accuracy on the training set rises from 64.0% to 80.5%, as expected after fine-tuning. More importantly, its test accuracy also rises—from 62.6% to 74.2%—an absolute gain of 11.6 points, putting it on par with the RL model’s 74.8%.
In summary, when distilling the base model with correct responses from the base model there is only minor improvement, whereas when distilling with correct responses from the RLVR-trained model there is significant improvement, on par with RLVR itself. This suggests that there is a qualitative difference in the two response types, and reveal the following insight: RLVR does more than merely increase the success probability for the less difficult questions—it enables the model to produce quality responses that were not present in its output distribution before training.
5.2 Qualitative Analysis
Figure 3: Qualitative analysis of Qwen2.5-1.5B-Math before and after RLVR training. (Left) Mean response length, and (Right) mean count of reflection keywords, grouped by correctness.
In addtion, given that RLVR improves the model’s ability to generate quality responses, we examine how RLVR changes the qualitative characteristics of responses. Prior work suggests that RLVR-trained models often produce longer answers and show reflective reasoning behaviors such as verification and backtracking DeepSeek-AI (2025); Gandhi et al. (2025); Liu et al. (2025); Yeo et al. (2025); Zeng et al. (2025). Following these, we compare responses from the base and RL models along two surface-level dimensions: response length and the frequency of reflection-related keywords (e.g., “let’s verify,” “alternatively,” “wait”).
As shown in Figure 3, in our small model settings (1.5B and 3B), we observe no significant difference in response length. Moreover, the RL model tends to produce more direct responses, using fewer reflection-related keywords (Full results in Appendix A.6). These patterns diverge from prior observations in the literature. These findings imply that surface-level traits, such as response length or the frequency of reflection keywords, may not reliably indicate response quality. It also underscores the need for developing better quality evaluation criteria in future work.
6 Why Does Teacher Distillation Typically Increase Capability when RLVR Keeps It Unchanged?
Distillation from teacher reasoning models is known to be another effective approach for improving accuracy Min et al. (2024); Muennighoff et al. (2025); Ye et al. (2025). Here, we take a step further and investigate: Can teacher distillation also improve capability, and under what conditions does such improvement occur?
6.1 Capability Analysis
Figure 4: Pass@$k$ comparisons across AIME25 and MATH500 datasets for the 1.5B model and their distillation-trained variants. For the MATH500 results, we show the results on 50 questions with the lowest success rates under the base model to better visualize the performance gaps between models.
Yue et al. briefly explored this issue by comparing two models: Qwen-2.5-Math-7B Yang et al. (2024) and DeepSeek-R1-Distill-Qwen-7B DeepSeek-AI (2025). The latter is a publicly available model obtained by distilling 800K DeepSeek-R1 responses into the Qwen-2.5-Math-7B student model. Their experiments showed that the distilled model demonstrates improved capability, as evidenced by higher pass@$k$ scores. However, it remains unclear why the distillation process led to the observed capability improvement. Teacher model responses contain two key components: (1) the model’s reasoning patterns, and (2) its domain knowledge. In the case of DeepSeek-R1-Distill-Qwen-7B, the student model was trained on up to 800K responses generated by DeepSeek-R1 DeepSeek-AI (2025), making it highly likely that new mathematical knowledge absent in the original student model’s pre-training data was also introduced during distillation. As a result, it is difficult to determine whether the observed gains in capability stem from adopting more effective reasoning pattern or from learning new knowledge.
To disentangle these effects, we conduct a comparative study using three models: (1) Qwen2.5-Math-1.5B, a non-reasoning base model; (2) DeepSeek-R1-Distill-Qwen-1.5B, which, like its 7B counterpart, was trained on 800K responses from DeepSeek-R1 DeepSeek-AI (2025) and likely benefits from both new knowledge and improved reasoning; (3) a distilled model we train ourselves, designed to isolate the effect of reasoning-pattern transfer without introducing any new domain knowledge beyond what the base model already possesses. For simplicity, we refer to these as the base model, the DeepSeek model, and the reasoning-only model.
We train the reasoning-only model using the following procedure. Starting with the base model, we generate 256 responses for each of the 7,500 questions in the MATH train set. We exclude any question for which the model achieves a zero success rate. Since the base model answers each remaining question correctly at least once, we treat these as in-distribution. As the teacher, we use QwQ-32B Qwen (2024), which we verified to have higher capability than the student model (see Appendix A.7 for details). For each in-distribution question, QwQ-32B generates 8 candidate responses. From these, we randomly select up to 4 correct ones (using all available if fewer than 4 exist) and conduct SFT on the base model with them. Distilling in this manner should not introduce any new knowledge beyond what the base model already has. The resulting distilled model achieves 68% accuracy on the MATH 500 dataset, outperforming the base model’s 60%, indicating a successful distillation.
We evaluate the capability of the three models by conducting pass@$k$ experiments on the AIME25 OpenCompass (2025) and MATH500 dataset. Results are shown in Figure 4. We observe the following patterns consistently across both datasets. (1) The DeepSeek model outperforms the base model across the entire range of $k$ values. For AIME25, it achieves a pass@256 score of 70.0%, compared to 56.7% for the base model. To ensure this gap does not close at higher $k$, we extend evaluation of the base model up to pass@512 and confirm that its performance plateaus, remaining at 56.7%. A similar pattern is observed in MATH500. This shows that the DeepSeek model achieves gains in both accuracy and capability, consistent with (Yue et al., 2025). (2) In contrast, the reasoning-only model initially outperforms the base model at lower $k$, but the curves converge and even cross as $k$ increases. We also replicate the experiment with Qwen2.5-3B and observe the same convergence pattern (See Appendix A.8 for results). These results show that the reasoning-only model improves accuracy but does not expand capability, similar to the case of RLVR discussed in Section 4.
These results lead to following insight: Teacher distillation improves capability only when it introduces new knowledge. Learning reasoning patterns alone boosts accuracy but not capability.
6.2 In-Depth Analysis Based on Question Difficulty
Figure 5: Change in success rates (absolute difference in %) for the DeepSeek model and the reasoning-only model, grouped by base model success rate bins.
To better understand the dynamics of the two types of distillation, we conduct an analysis similar to that in Section 4.2, comparing model performance across question bins grouped by base model success rates, based on 256 sampled responses. As shown in Figure 5, we observe the following patterns: (1) The DeepSeek model shows substantial improvement across all bins, including those with zero or near-zero success rates. (2) In contrast, the reasoning-only model shows improvement primarily in bins with moderately high success rates, but little gain in the zero or near-zero bins, mirroring the behavior of RLVR discussed in Section 4.
As in Section 4.2, we track how per-question success rates change before and after two different distillation settings and visualize the transitions in Figure 6. We observe: (1) For the DeepSeek model, questions consistently move to higher success-rate bins, even those that started in low-success bins. For instance, in the [1–4] bin, only 11.1% (2 out of 18) drop to the [0] bin, and in the [5–16] bin, no question moves downward. (2) In contrast, for the reasoning-only model, we interestingly observe the same "sacrificing difficult problems" effect seen in RLVR. In the [1–4] bin, 38.9% (7 out of 18) drop to the [0] bin, and in the [5–16] bin, 29.4% (10 out of 34) move to lower bins.
These results suggest the following insight: Distillation with new knowledge improves both accuracy and capability because it enables the model to solve questions across all difficulty levels, including the most challenging ones. In contrast, reasoning-only distillation improves accuracy but not capability because, like RLVR, it focuses on easier questions—often at the cost of performance on the hardest ones.
Figure 6: Success rate transition matrix showing redistribution of questions in two distillation settings on MATH 500 test set.
7 Conclusions
Recent prior work has shown that RLVR improves accuracy but not capability, whereas teacher distillation from a strong teacher typically improves both accuracy and capability. In this paper, we provide explanations for these two phenomena, which are backed up with extensive empirical experimentation. In summary, the RLVR training algorithm focuses on improving the accuracy of the easier problems while paying less attention to the difficult problems; thus the difficult problems on the whole continue to have very low success rates, and thus the model’s overall capability remains unchanged. Teacher distillation is typically done with a strong teacher model which includes knowledge not present in the original base model; this leads to an increase in capability. Comparing RLVR and teacher distillation is thus an apples-to-oranges comparison, in that RLVR only uses the resources it has on hand, whereas teacher distillation augments the model with new resources.
Our specific contributions are as follows:
  1. 1.
    We explain why RLVR improves accuracy but not capability, showing that it disproportionately favors easier questions to the detriment of the harder ones—often degrading performance for difficult questions.
  2. 2.
    We demonstrate that RLVR leads to better quality responses, as evidenced by self-distillation experiments, even though surface-level indicators such as response length or reflection keywords fail to capture this improvement.
  3. 3.
    We show that teacher distillation improves capability only when it introduces new knowledge, whereas distilling only reasoning patterns behaves similarly to RLVR—boosting accuracy at the expense of capability.
Together, these findings offer a clearer understanding of how RLVR and distillation shape reasoning behavior in LLMs.
8 Limitations
While our study offers an in-depth analysis of RLVR and distillation, several limitations point to future directions. First, our experiments are limited to relatively small models (1.5B and 3B), a single domain (mathematics), and a single RL algorithm family (GRPO & Dr. GRPO). Larger models, alternative domains, or different RL algorithms may exhibit different dynamics. A more comprehensive study across model scales, tasks, and training methods is needed to test the generality of our findings. Second, while we observe the importance of response quality in distillation across different experiments, the underlying factors that determine such quality remain underexplored. Future research may investigate what makes a response effective for improving accuracy and generalization, and propose more principled metrics for evaluating reasoning quality. Third, our results suggest that capability gains in distillation primarily stem from the introduction of new knowledge absent from the base model. While our approach offers a practical way to isolate this effect, future work could develop more rigorous frameworks to better characterize, quantify, and control knowledge transfer during distillation.
Appendix A Appendix
A.1 Accuracy vs. Capability Example
As discussed in Section 3, we provide an example to illustrate that a model can have higher accuracy but lower capability on an evaluation dataset with more than one question.
Recall the definitions:
$\mathrm{Acc}(M)$ $=\frac{1}{N}\sum_{i=1}^{N}p_{i}^{M},$
$\mathrm{Cap}_{k}(M)$ $=\frac{1}{N}\sum_{i=1}^{N}\left(1-(1-p_{i}^{M})^{k}\right).$
We compare two models, $M_{1}$ and $M_{2}$, on a toy dataset of $N=3$ questions. Their single-attempt success probabilities $p_{i}^{M}$ are shown below:
Question $p_{i}^{M_{1}}$ $p_{i}^{M_{2}}$
1 0.9 0.5
2 0.9 0.5
3 0.003 0.5
Table 2: Single-pass success probabilities for models $M_{1}$ and $M_{2}$.
We first compute the accuracy of two models on this toy dataset.
$\mathrm{Acc}(M_{1})$ $=\frac{1}{3}(0.9+0.9+0.003)=0.601,$
$\mathrm{Acc}(M_{2})$ $=\frac{1}{3}(0.5+0.5+0.5)=0.5.$
Thus, $M_{1}$ has higher accuracy.
We now compute capability with $k=256$, which is large enough to expose the low success probability on Question 3 for $M_{1}$:
Using the formula:
$p_{i,k}^{M}=1-(1-p_{i}^{M})^{k},$
we compute:
Model $M_{1}$:
$p_{1,256}^{M_{1}}$ $=1-(1-0.9)^{256}\approx 1,$
$p_{2,256}^{M_{1}}$ $=1-(1-0.9)^{256}\approx 1,$
$p_{3,256}^{M_{1}}$ $=1-(1-0.003)^{256}\approx 0.537.$
$\mathrm{Cap}_{256}(M_{1})$ $=\frac{1}{3}(1+1+0.537)\approx 0.845.$
Model $M_{2}$:
$p_{i,256}^{M_{2}}$ $=1-(1-0.5)^{256}=1-2^{-256}\approx 1\quad\text{for all }i,$
$\mathrm{Cap}_{256}(M_{2})$ $=\frac{1}{3}(1+1+1)=1.0.$
As shown, although $M_{1}$ has significantly higher probabilities to the first two questions—resulting in higher overall accuracy—its probability on the third question is extremely low. As a result, even with many sampling attempts, $M_{1}$ is unlikely to solve all questions. In contrast, $M_{2}$ maintains moderate but consistent success probabilities across all three questions, which leads to a higher chance of solving every question at least once when given sufficient attempts.
A.2 Pass@$k$ Experiments Results Before & After RLVR
In this paper, we used two models to evaluate the effect of RLVR training: Qwen2.5-1.5B-Math, and Qwen2.5-3B. For corresponding RL model of 1.5B model, we used the Qwen2.5-Math-1.5B-Oat-Zero, a publicly available model trained with MATH train dataset by Liu et al.. For Qwen2.5-3B, we conducted the RLVR training ourselves, also with MATH train dataset. Further details for training can be found at Appendix A.9 and A.10, respectively.
Split Model Qwen2.5-1.5B-Math Qwen2.5-3B
Accuracy Maj@256 Pass@256 Accuracy Maj@256 Pass@256
Train Base 64.0% 76.8% 97.2% 59.3% 80.9% 92.7%
RL 80.9% 82.1% 97.1% 67.9% 82.2% 92.1%
Difference +16.9% +5.3% -0.1% +8.6% +1.3% -0.6%
Test Base 60.6% 72.0% 97.2% 54.9% 76.5% 95.8%
RL 74.2% 80.8% 97.0% 63.6% 79.5% 95.8%
Difference +13.9% +8.8% -0.2% +8.7% +3.0% +0.0%
Table 3: Performance comparison of base and RL models for Qwen2.5-1.5B-Math and Qwen2.5-3B
Figure 7: Pass@k comparison between base and RLVR-trained models on train and test sets.
Similar to the work done by Yue et al., we conducted the pass@$k$ experiments with these models. For both the base and RL models, we generated 256 responses per question on the MATH train set and MATH500 test set. Using these responses, we estimated accuracy and pass@$k$ capability for $k=1$ to 256, following the metric defined in Section 3.2. Additionally, we computed majority vote accuracy (maj@256), which is the percentage of questions where the most frequent answer among the 256 responses is correct.
As expected, we observed that RLVR significantly improved both accuracy and majority vote performance across training and test sets. As shown in Table 3, these gains appeared consistently in both the 3B and 1.5B models, indicating that RLVR leads to generalizable improvement in accuracy without signs of overfitting. In contrast, we observed no meaningful improvement in capability. For both the 1.5B and the 3B models, pass@k either remained stable or slightly declined across the training and test sets. As shown in Figure 7, the RL model outperformed the base model at small $k$, but their curves converged as $k$ increases—a pattern consistent with prior work Shao et al. (2024); Yue et al. (2025).
A.3 Question-Difficulty-Based Analysis Results
As discussed in Sections 4, we performed detailed analyses based on question difficulty across different training settings. The results are presented below. Figure 8 shows success rate improvements across difficulty bins for both 1.5B and 3B models on train and test. Figure 9 presents the corresponding transition matrices that illustrate how questions move between success rate bins before and after training.
(a) Qwen2.5-1.5B-Math (Train)
(b) Qwen2.5-1.5B-Math (Test)
(c) Qwen2.5-3B (Train)
(d) Qwen2.5-3B (Test)
Figure 8: Change in success rates (absolute %) across difficulty bins for Qwen2.5-1.5B-Math and Qwen2.5-3B on the MATH training and test sets. In both models, RLVR significantly improves questions in the mid-success bins (e.g., [17–64], [65–128]), but yields minimal gains in the lowest bins ([0], [1–4]).
(a) Qwen2.5-1.5B-Math (Train)
(b) Qwen2.5-1.5B-Math (Test)
(c) Qwen2.5-3B (Train)
(d) Qwen2.5-3B (Test)
Figure 9: Transition matrices comparing base and RLVR success-rate bins for Qwen2.5-1.5B-Math and Qwen2.5-3B. Each cell shows the percentage and count of questions moving between success bins. Most upward transitions occur from mid-success bins; questions in low-success bins are more likely to remain unchanged or regress.
A.4 Entropy Analysis
(a) Qwen2.5-1.5B-Math
(b) Qwen2.5-3B
Figure 10: Comparison of output entropy across base and RLVR-trained models, computed over 256 responses per question. Entropy is measured by grouping responses with identical final answers.
As discussed in Section 4.2, we performed an entropy analysis to examine whether RLVR reduces response diversity and causes the model to concentrated on fewer outputs. For both the base and RL models, we computed the entropy of the answers in the 256 responses for each question and then average the entropy values over the questions. Entropy values were measured by grouping responses that yield the same final answer, regardless of correctness. We conducted the experiment for both Qwen2.5-1.5B-Math and Qwen2.5-3B.
As shown in Figure 10, output entropy dropped noticeably after RLVR. More importantly, this reduction consistently appeared across all difficulty levels, including questions where the base model had zero or near-zero success rate. These results support the hypothesis that RLVR reinforces greater consistency, but on harder questions, this often leads the model to focus on incorrect responses—making it less likely to recover even occasional correct answers.
A.5 Self-Distillation Results
Table 4 reports the full accuracy results for the self-distillation experiments introduced in Section 5.1. All fine-tuning used the same hyperparameter configuration (Appendix A.11).
Model Student Model Teacher Model Train Accuracy Test Accuracy
Base model 64.0% 62.6%
Base model Base model 74.7% (+10.7%) 63.4% (+0.8%)
Qwen2.5-1.5B-Math RL model 80.9% 74.8%
RL model RL model 84.4% (+3.5%) 74.4% (-0.4%)
Base model RL model 80.5% (+16.5%) 74.2% (+11.6%)
Base model 59.3% 54.9%
Base model Base model 73.6% (+14.3%) 58.7% (+3.8%)
Qwen2.5-3B RL model 67.9% 63.6%
RL model RL model 72.1% (+4.2%) 64.4% (+0.8%)
Base model RL model 73.6% (+14.3%) 64.5% (+9.6%)
Table 4: Self-distillation results for Qwen2.5-1.5B-Math and Qwen2.5-3B across different student–teacher configurations. Accuracy values in parentheses reflect improvements over the corresponding student model before fine-tuning.
A.6 Qualitative Analysis of Responses Before & After RLVR
In Section 5.2, we compared responses from Qwen2.5-1.5B-Math and Qwen2.5-3B before and after RLVR training along two dimensions: response length and the use of reflection-related keywords (e.g., "let’s verify", "alternatively", "wait"). Here, we present the complete results.
A.6.1 Response Length
(a) Qwen2.5-1.5B-Math: Mean response length grouped by correctness.
(b) Qwen2.5-1.5B-Math: Mean response length by difficulty and correctness.
(c) Qwen2.5-3B: Mean response length grouped by correctness.
(d) Qwen2.5-3B: Mean response length by difficulty and correctness.
Figure 11: Comparison of response lengths between base and RLVR-trained models across model sizes and difficulty levels. Top row: Qwen2.5-1.5B-Math; bottom row: Qwen2.5-3B. Left: Mean response length grouped by correctness. Right: Mean response length further stratified by difficulty. In both models, RLVR did not increase response length, and correct responses tended to be more concise.
For both 1.5B and 3B models, we generated 256 responses per MATH500 question from both the base and RL models and computed mean response lengths. We also separated responses by correctness. As shown in Figure 11, there was no substantial difference in average length between the two models. In both cases, correct responses were consistently shorter than incorrect ones.
To control for the correlation between question difficulty and correctness, we grouped questions into four bins based on how many of the 256 base model responses were correct—higher bins indicating easier questions. Within each bin, we compared mean response lengths by correctness. As shown in the figure, both models exhibited the same trend: correct responses were consistently shorter, and overall response lengths remained similar, indicating that RLVR did not increase response length.
A.6.2 Reflection-Related Keywords
(a) Qwen2.5-1.5B-Math: Mean count of reflection keywords grouped by correctness.
(b) Qwen2.5-1.5B-Math: Reflection keyword frequency by difficulty and correctness.
(c) Qwen2.5-3B: Mean count of reflection keywords grouped by correctness.
(d) Qwen2.5-3B: Reflection keyword frequency by difficulty and correctness.
Figure 12: Reflection-related keyword analysis across base and RLVR-trained models. Top row: Qwen2.5-1.5B-Math; bottom row: Qwen2.5-3B. Left: Mean count of reflection-related keywords, grouped by correctness. Right: Keyword frequency stratified by question difficulty and correctness. Across both models, RLVR-trained responses consistently contain fewer reflective phrases.
Prior work suggests that RLVR elicits more non-linear reasoning in model outputs DeepSeek-AI (2025); Gandhi et al. (2025). To test this, we analyzed the presence of predefined reflection-related phrases. The full list of the phrases are available in Table 5.
Reflection-Related Keywords
actually aha alternatively
another approach checking our work correction:
different method double-check hmm
however I made a mistake I need to reconsider
I realize let me recalculate let me think
let’s check let’s reconsider looking back
make sure ok on second thought
retracing to be sure to confirm
verify wait we could also
Table 5: List of reflection-related phrases used for qualitative analysis.
Using the same setup as the response length analysis, we examined 256 responses per question across all MATH500 test questions for both base and RL models. For each response, we counted occurrences of reflection keywords and stratified results by correctness and difficulty.
As shown in Figure 12, the RL model exhibited substantially fewer reflection keywords than the base model. The figure further shows that, while the base model showed little variation across correctness levels, the RL model consistently used fewer reflection phrases in correct answers across all difficulty bins. These responses were generally more direct and less exploratory.
A.7 QwQ-32B Capability Experiment
Figure 13: Pass@$k$ results of QwQ-32B, Qwen2.5-3B-Math, and Qwen2.5-1.5B-Math on AIME 25
In Section 6, we selected QwQ-32B as the teacher model for our reasoning-only distillation experiment. To ensure a fair test of whether distillation can improve capability without introducing new knowledge, the teacher must have higher capability than the student models—Qwen2.5-3B and Qwen2.5-1.5B-Math.
To validate this, we conducted a pass@$k$ evaluation on AIME 25 using 64 responses per question from QwQ-32B, and compared the results with the two student models. As shown in Figure 13, QwQ-32B consistently outperforms both students across all $k$ values, with no sign of convergence. Notably, its pass@64 score reached 76.7%, compared to just 43.3% and 56.7% at pass@256 for Qwen2.5-3B and Qwen2.5-1.5B-Math, respectively. These results confirm that QwQ-32B has substantially higher capability, making it a suitable teacher model for our distillation setup.
A.8 Teacher Distillation Pass@$k$ results
As discussed in Section 6, we conducted the pass@$k$ on AIME 25 and MATH 500 for both Qwen2.5-1.5B-Math and Qwen2.5-3B and each their 2 distilled variants. The results are shown below in Figure 14.
Figure 14: Pass@$k$ comparisons across AIME25 and MATH500 datasets for both 1.5B (top) and 3B (bottom) models and their distillation-trained variants. For the MATH500 results of the 1.5B models, we show performance on the 50 questions with the lowest base-model success rates to better highlight the differences.
A.9 Qwen2.5-Math-1.5B Training Details
In this paper, we used two models as base models: Qwen2.5-1.5B-Math and Qwen2.5-3B. For the RLVR-trained version of the 1.5B model, we used Qwen2.5-Math-1.5B-Oat-Zero [^2] https://huggingface.co/sail/Qwen2.5-Math-1.5B-Oat-Zero, a publicly available model trained by Liu et al.. According to their report, the model was trained with Dr.GRPO Liu et al. (2025), a variant of the GRPO algorithm Shao et al. (2024) designed to remove response length and question difficulty biases. The model was trained on questions from level 3 to 5 from the MATH training set. For the 3B model, we performed RLVR training ourselves. Training details are shown right below in Appendix A.10
A.10 Qwen2.5-3B RLVR Training Details
For RLVR training of Qwen2.5-3B, we used the GRPOTrainer from the TRL [^3] https://github.com/huggingface/trl library, which implements the standard GRPO algorithm. The model was trained on the full MATH training set, consisting of 7,500 questions.
A.10.1 Prompt Setting
Prior work has shown that the performance of smaller models can be sensitive to prompt design Hochlehnert et al. (2025); Liu et al. (2025). Following Liu et al., we evaluated three prompt formats, as listed below. We ultimately adopted Template 3 (question only), which yielded the best performance.
[…]
A.10.2 Reward Function
We adopted a minimalistic reward setting. A response received a reward of 1 if it contained the correct final answer, and -1 otherwise. Answer verification was performed using the math_verify [^4] https://github.com/huggingface/Math-Verify package.
$R(q,a,r)=\begin{cases}1&\text{if the response $r$ to question $q$ matches the ground truth answer $a$}\\ -1&\text{otherwise}\end{cases}$
A.10.3 RLVR Training Hyperparameters
Table 6 summarizes the key hyperparameters used in RLVR training for the Qwen2.5-3B model.
Hyperparameter Value
Optimizer AdamW
Learning rate scheduler Constant
Maximum token length 4000
Temperature 0.9
Top-$p$ 1.0
Top-$k$ 50
Number of generations (per question) 10
Global batch size 4 (per device) $\times$ 7 (GPUs) $\times$ 10 (accumulation) = 280
Learning rate $1\times 10^{-6}$
Gradient clipping (max grad norm) 0.1
Number of gradient steps 225
Warmup steps 20
Mixed precision bf16
Table 6: Key hyperparameters used for RLVR training of Qwen2.5-3B.
A.10.4 Training Progress and Evaluation
Figure 15: Change in response length (token counts) over training and accuracy on MATH 500 across checkpoints.
During RLVR training, we evaluated the model every 25 gradient steps. To ensure statistical robustness, we followed the recommendation of Hochlehnert et al. Hochlehnert et al. (2025) and sampled responses 10 times per checkpoint, reporting the mean accuracy. For each evaluation, we used a temperature of 0.8 and a top-$p$ of 0.9.
As shown in Figure 15, accuracy peaked at step 150, reaching 63.6%, and then plateaued. We selected this checkpoint as the RL model used throughout our experiments. The figure also shows the average response length over training. As discussed in Section 5.2, response length remained stable, showing no significant growth.
A.11 Distillation Training Hyperparameters
For all the self-distillation experiments in Section 5.1 and teacher distillation in Section 6, we used the supervised fine-tuning (SFT) hyperparameters listed in Table 7.
Hyperparameter Value
Optimizer AdamW
Learning rate scheduler Constant
Weight decay $1\times 10^{-4}$
Warmup steps 25
Max sequence length 32,768
Global batch size 4
Mixed precision bf16
Table 7: Key hyperparameters used for supervised fine-tuning in distillation experiments.
A.12 Response Generation Details
We used vLLM [^5] https://docs.vllm.ai library Kwon et al. (2023) for response generation and math_verify [^6] https://github.com/huggingface/Math-Verify package for response grading.
We used temperature 0.9, top-$p$ of 1.0, and top-$k$ of 50 for all models, except where noted below. These settings were chosen to ensure response diversity. Unless otherwise specified, we used the question-only template (Template 3).
For Qwen2.5-Math-1.5B-Oat-Zero, we used the same sampling hyperparameters but followed the Qwen prompt format (Template 2), as recommended in the user guideline. [^7] https://huggingface.co/sail/Qwen2.5-Math-1.5B-Oat-Zero
For QwQ-32B, we used temperature 0.6, top-$p$ 0.95, and top-$k$ 50. We followed the R1 prompt template (Template 1), as recommended in the user guideline. [^8] https://huggingface.co/Qwen/QwQ-32B#usage-guidelines
For DeepSeek-R1-Distill-Qwen-1.5B, we used temperature 0.6, top-$p$ 0.95, and top-$k$ 50. We followed the R1 prompt template (Template 1), as recommended in the user guideline. [^9] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-1.5B#usage-recommendations